Term after term find the first five terms of the given general term of an=2n-1
Answer by farohw(175)


Let n = 1, 2, 3, 4, 5

For n = 1: a_n = 2n-1 ---> a_1 = 2(1) - 1 = 1

For n = 2: a_2 = 2(2) - 1 = 3

For n = 3: a_3 = 2(3) - 1 = 5

For n = 4: a_4 = 2(4) - 1 = 7

For n = 5: a_5 = 2(5) - 1 = 9

The first five terms of the sequence is 1, 3, 5, 7, 9.